# Primordial gravitational waves in Wheeler-DeWitt non-commutative linearized branch-cut quantum gravity

### 12 Jan 2024 v1

@@ III Non-commutative Branch-cut Quantum Gravity
$\xi=\varphi(u,v),\quad\mbox{and}\quad\eta=\eta(u,v)\,,$ (19)
where $\eta(u,v)$ is a differentiable function that satisfies the non-degeneracy condition of the Jacobian $D(\xi,\eta)/D(x,y)$ in the given domain and $\varphi(u,v)$ represents the general integral. As a result, equation (11) is reduced to the canonical form
${\cal H}(\xi,\eta)\Psi(\xi,\eta)=\frac{1}{2}\frac{N}{\eta}\Biggl{[}\frac{\partial^{2}}{\partial\eta^{2}}+\frac{\gamma}{\eta^{3\alpha-1}}\frac{\partial}{\partial\eta}+g_{r}-g_{m}\eta-g_{k}\eta^{2}-g_{q}\eta^{3}+g_{\Lambda}\eta^{4}+\frac{g_{s}}{\eta^{2}}+\frac{\alpha}{\eta^{3\alpha-2}}-\frac{\alpha\xi}{\eta^{3\alpha-1}}-\frac{i}{\eta^{3\alpha-1}}\frac{\partial}{\partial\xi}\Biggr{]}\Psi(\xi,\eta)=0\,,$ (20)
@@
${\cal H}(\xi,\eta)\Psi(\xi,\eta)=\frac{1}{2}\frac{N}{\eta}\Biggl{[}-p^{2}_{\eta,\gamma,\alpha}+g_{r}-g_{m}\eta-g_{k}\eta^{2}-g_{q}\eta^{3}+g_{\Lambda}\eta^{4}+\frac{g_{s}}{\eta^{2}}+\frac{\alpha}{\eta^{3\alpha-2}}-\frac{\alpha\xi}{\eta^{3\alpha-1}}+\frac{1}{\eta^{3\alpha-1}}p_{\xi}\Biggr{]}\Psi(\xi,\eta)=0\,.$ (21)
@@
$p_{\eta,\gamma,\alpha}\equiv-\frac{i|\gamma|}{\eta^{3\alpha-1}}\frac{\partial}{\partial\eta}\mapsto\frac{\gamma}{\eta^{3\alpha-1}}\Bigl{(}-i\frac{\partial}{\partial\eta}\Big{)}=\frac{\gamma}{\eta^{3\alpha-1}}p_{\eta}\,.$ (23)
With the particular choice $\alpha=1/3$, that allows a separation of variables, with $\Psi(\xi,\eta)=\Psi(\xi)\Psi(\eta)$, equation (21), reduces to the following equations [9]
$\Biggl{(}\frac{\partial^{2}}{\partial\eta^{2}}+\gamma\frac{\partial}{\partial\eta}+V(\eta)\Biggr{)}\Psi(\eta)=0\,,$ (24)
@@
$V(\eta)=\tilde{g}_{r}-\tilde{g}_{m}\eta-g_{k}\eta^{2}-g_{q}\eta^{3}+g_{\Lambda}\eta^{4}+\frac{g_{s}}{\eta^{2}}\,,$ (25)
@@
$\Biggl{(}i\frac{\partial}{\partial\xi}-\frac{1}{3}\xi+{\cal C}\Biggr{)}\Psi(\xi)=0\,;$ (26)
@@
$\Psi(\xi)=i\Biggl{(}{\cal C}\xi-\frac{1}{3}\xi^{2}\Biggr{)}\,.$ (27)
III.1 Probability interpretation of the wave function of the Universe
@@
$\Bigl{(}\partial_{\mu}\partial^{\mu}+m^{2}\Bigr{)}\Phi(x)=0\,,$ (28)
@@
$j=\frac{i}{2}\Bigl{(}\Phi^{*}\,\nabla\cdot\Phi-\Phi\,\nabla\cdot\Phi^{*}\Bigr{)},$ (29)
@@
$j^{\mu}(x)=i\Bigl{[}\Phi^{*}(x)\partial^{\mu}\Phi(x)-\Bigl{(}\partial^{\mu}\Phi^{*}(x)\Bigr{)}\Phi(x)\Bigr{]}.$ (30)
Combining (28) and (30), we obtain $\partial_{\mu}j^{\mu}(x)=0$, so the four-current $j^{\mu}(x)$ is a conserved quantity. However, the density
$\rho(x)=i\Bigl{[}\Phi^{*}(x)\frac{\partial\Phi(x)}{\partial t}-\frac{\partial\Phi^{*}(x)}{\partial t}\Phi(x)\Bigr{]},$ (31)
@@
$\langle\Phi_{1}(x)|\Phi_{2}(x)\rangle=\frac{i}{2}\int_{{\cal R}^{3}}d^{3}x\rho(x)_{\Phi_{1}\Phi_{2}}\,,$ (32)
@@
$\rho(x)_{\Phi_{1}\Phi_{2}}=\Bigl{[}\Phi_{2}(x)\frac{\partial\Phi^{*}_{1}(x)}{\partial t}-\frac{\partial\Phi^{*}_{2}(x)}{\partial t}\Phi_{1}(x)\Bigr{]}\,,$ (33)
@@
$i\frac{\partial}{\partial t}\Psi=\hat{H}\Psi\,.$ (35)
@@
${\cal J}=\frac{i}{2}\Bigl{(}\Psi^{\dagger}\,\nabla\cdot\Psi-\Psi\,\nabla\cdot\Psi^{\dagger}\Bigr{)};\,\,\,\mbox{with}\,\,\nabla\cdot{\cal J}=0\,.$ (36)
@@
For Rovelli [46] the absence of time is a feature of the classical Hamilton-Jacobi formulation of general relativity, and the wave function is only a function of the “3-geometry”, namely the equivalence class of metrics under a diffeomorphism, and not of the specific coordinate dependent form of the metric tensor. According to the second law of thermodynamics, forward in time represents the direction in which entropy increases and in which we obtain information, so the flow of time would represent a subjective feature of the universe, not an objective part of physical reality [46]. In this realm, in which the observable universe does not exhibit time-reversal symmetry, events, rather than particles or fields, are the basic constituents of the universe, implying that the evolution of physical quantities is related to the description of the relationship between events [44, 45, 46]. For instance, given the wave function of the universe as a functional constrained to a region configuration of a super-space that contains a three-surface $\Sigma$ and matter fields configuration, represented by $\phi$, where the metric is described by $h_{ij}$ the corresponding WdW wave function $\Psi(h_{ij},\phi)$ may be interpreted according to [46], as stressed before, as describing the evolution of $\Psi(\phi)$ in the physical variable $\phi$. In the following we assume the proposal made by Hawking [57] for which $|\Psi[h_{ij},\phi,\Sigma]|^{2}$ is proportional to the probability $P({\cal A})$ of finding a 3-surface $\Sigma$ with metric $h_{ij}$ and matter field configuration $\phi$:
$*P({\cal A})\propto\int_{\cal A}|\Psi[h_{ij},\phi,\Sigma]|^{2}\,{}^{*}1$ (37)
@@
where $*{}^{*}$ is the Hoge dual in the metric $\Gamma(N)$, i.e. $*{}^{*}1$ is the volume element [57]. $\Gamma(N)$ in turn represents the superspace metric which does not depend linearly on $N$. In what follows, we adopt similar interpretations to Hawking’s prescription for $|\Psi(\eta)|^{2}$ and $|\Psi(\xi)|^{2}$, and to establish a connection with standard quantum mechanics we use the denomination “probability density”.
@@
Briefly, as highlighted by Hartle [62], in standard quantum mechanics, the probabilities associated with wave functions are represented by squares of amplitudes, and additionally, a criterion is needed to specify which sets of histories can have probabilities consistently assigned to them. And yet, as highlighted by Hartle, in standard quantum mechanics that criterion is measurement – probabilities can be consistently assigned to histories of measured alternatives and usually not otherwise. Hartle then recalled that a criterion based on measurements or observers cannot be fundamental in a quantum theory that seeks to explain the early universe where neither existed. A more general criterion for closed systems assigns probabilities to just those sets of histories for which there is vanishing interference between its individual members as a consequence of the system’s initial quantum state. Such sets of histories are said to be decohere. Decoherent sets of histories are what may be used for prediction and retrodiction in quantum cosmology for they may be assigned probabilities [62]. And this type of identification, in our view, with decoherent sets of histories, is in perfect harmony with Hawking’s prescription and with the modes of interpretation adopted by us for the dynamic evolutionary processes of the Universe’s wave function and the scale factor in branch-cut gravity.
III.2 The wave function of the Universe in the non-commutative approach
@@
In what follows, we use the powerful iterative Range-Kutta-Fehlberg numerical analysis for solving differential equations to find solutions for equation (24), without introducing numerical approximations, in order to describes the evolution of the wave function of the Universe, $\Psi(\eta)$. The boundary conditions of the solutions are based on the Bekenstein criterion, which provides an upper limit for the Universe’s entropy, following the proposition presented in [8, 9]. The total entropy of a black hole, according to the Bekenstein limit, is proportional to the number of Planck areas needed to cover the event horizon, where each area corresponds to one unit of entropy. In non-commutative branched gravitation, we assume that the primordial singularity is equally covered by a certain number of Planck areas, whose numerical value in turn corresponds to the total primordial entropy of the Universe. We assume that the dimensions of this boundary region correspond to the farthest points observable while still respecting causality. For this, we consider an appropriate distance, denoted as $d(t)$, between a pair of objects at any given moment $t$, and the corresponding distance, denoted as $d(t_{0})$, at a reference time $t_{0}$. We establish this relationship as $d(t)=|\eta(t)|d(t_{0})$. This means that the relationship between the two distances is modulated by the scale factor of the branch-cut Universe. This implies that for $t=t_{0}$ we have $|\eta(t_{0})|=1$. From a quantum probabilistic point of view, this condition implies a maximum probability of observation, $|\Psi(1)|^{2}=1$, assuming a normalized wave function. Thus, the boundary conditions considered in this contribution are, in the contraction sector $\Psi(-1)=-1$, while in the expansion region, $\Psi(1)=1$.
@@
Figures 1 and 2 depict the behavior of the solutions to equation (24), which pertains to the wave function of the Universe $\Psi(\eta)$ within a non-commutative approach. These results align with the predictions of branch-cut gravity , suggesting that the present Universe did not originate from nothing (see refs. [6, 7, 66]) or from a quantum loop [67]. Instead, it appears to have emerges from a prior phase before the current expansion phase. The implications of the non-commutative algebra on the acceleration of the Universe are most pronounced in figure 2, indicating the presence of a dynamic acceleration of the Universe driven by a force, whose work can be expressed as $W=-p_{f}dV$, where $p_{f}$ represents the strength of the pressure in the expanding region.
Figure 1: Typical sample solutions of the wave function of the Universe $\Psi(\eta)$ utilizing the non-commutative approach given by equation (24). The solutions are for initial conditions $\Psi(1)$ and $\Psi^{\prime}(1)$ and different values of the parameter $\gamma$. The parameter $\alpha$ was fixed at $1/3$. The parameter $\chi$ is implicitly embedded in the intrinsic structure of the scale factor $\eta$ as well as in the dual complementary scale factor $\xi$. The values of the running coupling constants and parameters are: $\tilde{g}_{r}=0.4$; $\tilde{g}_{m}=0.6185$; $g_{k}=1$; $g_{q}=0.7$; $g_{\Lambda}=0.333$; $g_{s}=0.03$; $\alpha=1/3$; ${\cal C}=1$.
@@
Figure 2: Typical solutions of the wave function of the Universe $\Psi(\eta)$ using the non-commutative approach given by equation (24). The parameter $\alpha$ was fixed at 1/3. The parameter $\chi$ is implicitly embedded in the intrinsic structure of the scale factor $\eta$ as well as in the dual complementary scale factor $\xi$. On the left side, in figures ((a), (c), and (e)) the values of the running coupling constants and parameters are as follows: $\tilde{g}_{r}=0.4$; $\tilde{g}_{m}=0.6185$; $g_{k}=1$; $g_{q}=0.7$; $g_{\Lambda}=0.333$; $g_{s}=-0.03$; $\alpha=1/3$; $\gamma=1$; ${\cal C}=1$. The boundary condition is set as $\Psi(-1)=-1$. On the right side, in figures ((b), (d), and (f)) , the values of the running coupling constants and parameters are: $\tilde{g}_{r}=0.4$; $\tilde{g}_{m}=0.6185$; $g_{k}=1$; $g_{q}=0.7$; $g_{\Lambda}=0.333$; $g_{s}=-0.03$; $\alpha=1/3$; $\gamma=-1$; ${\cal C}=1$. The boundary condition is set as $\Psi(1)=1$.
Another explanation could be based on the contribution of a new kind of dynamical energy fluid or field, that fills all space, whose effect on the expansion of the universe would be the opposite of that of “normal” matter and energy, as for instance the “quintessence” and “stiff” contributions, parameterized by the terms corresponding to the running coupling constants $g_{q}$ and $g_{s}$. Our results indicate that the term corresponding to stiff matter has expressive implications on the conformation of the potentials adopted in the present formulation, particularly with regard to the transition regions between the contraction and expansion phases of the branch-cut Universe.
Finally, when we compare the Universe expansion results corresponding to the commutative and non-commutative formulations, the configuration of matter, represented by the potentials (50), (54), (63), and (65), show expressive differences demonstrating that the non-commutative formulation imprints on the structure of spacetime a non-symmetrical redistribution of matter which captures, as stressed before, the short-distance scales.
@@
Figure 3: On the left side we show the complex plane solution of the wave function $\Psi(\xi)$, which is the dual complementary to the wave function of the Universe $\Psi(\eta)$ when utilizing the commutative approach given by equation (26). On the right side, we show the probability density of the wave function $\Psi(\xi)$ using Hawking’s prescription given by equation (37).
III.3 Dual spaces in quantum mechanics

### 03 Feb 2024 v2

@@ 3 Non-commutative Branch-cut Quantum Gravity
$\xi=\varphi(u,v),\quad\mbox{and}\quad\eta=\eta(u,v)\,,$ (3.18)
where $\eta(u,v)$ is a differentiable function that satisfies the non-degeneracy condition of the Jacobian $D(\xi,\eta)/D(x,y)$ in the given domain and $\varphi(u,v)$ represents the general integral. As a result, equation (3.10) is reduced to the canonical form
${\cal H}(\xi,\eta)\Psi(\xi,\eta)$ $=$ $\frac{1}{2}\frac{N}{\eta}\Biggl{[}\frac{\partial^{2}}{\partial\eta^{2}}+\frac{\gamma}{\eta^{3\alpha-1}}\frac{\partial}{\partial\eta}+g_{r}-g_{m}\eta-g_{k}\eta^{2}-g_{q}\eta^{3}+g_{\Lambda}\eta^{4}$ (3.19)
$+\frac{g_{s}}{\eta^{2}}+\frac{\alpha}{\eta^{3\alpha-2}}-\frac{\alpha\xi}{\eta^{3\alpha-1}}-\frac{i}{\eta^{3\alpha-1}}\frac{\partial}{\partial\xi}\Biggr{]}\Psi(\xi,\eta)=0\,,$
@@
${\cal H}(\xi,\eta)\Psi(\xi,\eta)$ $=$ $\frac{1}{2}\frac{N}{\eta}\Biggl{[}-p^{2}_{\eta,\gamma,\alpha}+g_{r}-g_{m}\eta-g_{k}\eta^{2}-g_{q}\eta^{3}+g_{\Lambda}\eta^{4}$ (3.20)
$+\frac{g_{s}}{\eta^{2}}+\frac{\alpha}{\eta^{3\alpha-2}}-\frac{\alpha\xi}{\eta^{3\alpha-1}}+\frac{1}{\eta^{3\alpha-1}}p_{\xi}\Biggr{]}\Psi(\xi,\eta)=0\,.$
@@
$p_{\eta,\gamma,\alpha}\equiv-\frac{i|\gamma|}{\eta^{3\alpha-1}}\frac{\partial}{\partial\eta}\mapsto\frac{\gamma}{\eta^{3\alpha-1}}\Bigl{(}-i\frac{\partial}{\partial\eta}\Big{)}=\frac{\gamma}{\eta^{3\alpha-1}}p_{\eta}\,.$ (3.22)
With the particular choice $\alpha=1/3$, that allows a separation of variables, with $\Psi(\xi,\eta)=\Psi(\xi)\Psi(\eta)$, equation (3.20), reduces to the following equations [9]
$\Biggl{(}\frac{\partial^{2}}{\partial\eta^{2}}+\gamma\frac{\partial}{\partial\eta}+V(\eta)\Biggr{)}\Psi(\eta)=0\,,$ (3.23)
@@
$V(\eta)=\tilde{g}_{r}-\tilde{g}_{m}\eta-g_{k}\eta^{2}-g_{q}\eta^{3}+g_{\Lambda}\eta^{4}+\frac{g_{s}}{\eta^{2}}\,,$ (3.24)
@@
$\Biggl{(}i\frac{\partial}{\partial\xi}-\frac{1}{3}\xi+{\cal C}\Biggr{)}\Psi(\xi)=0\,;$ (3.25)
@@
$\Psi(\xi)=i\Biggl{(}{\cal C}\xi-\frac{1}{3}\xi^{2}\Biggr{)}\,.$ (3.26)
3.1 Probability interpretation of the wave function of the Universe
@@
$\Bigl{(}\partial_{\mu}\partial^{\mu}+m^{2}\Bigr{)}\Phi(x)=0\,,$ (3.27)
@@
$j=\frac{i}{2}\Bigl{(}\Phi^{*}\,\nabla\cdot\Phi-\Phi\,\nabla\cdot\Phi^{*}\Bigr{)},$ (3.28)
@@
$j^{\mu}(x)=i\Bigl{[}\Phi^{*}(x)\partial^{\mu}\Phi(x)-\Bigl{(}\partial^{\mu}\Phi^{*}(x)\Bigr{)}\Phi(x)\Bigr{]}.$ (3.29)
Combining (3.27) and (3.29), we obtain $\partial_{\mu}j^{\mu}(x)=0$, so the four-current $j^{\mu}(x)$ is a conserved quantity. However, the density
$\rho(x)=i\Bigl{[}\Phi^{*}(x)\frac{\partial\Phi(x)}{\partial t}-\frac{\partial\Phi^{*}(x)}{\partial t}\Phi(x)\Bigr{]},$ (3.30)
@@
$\langle\Phi_{1}(x)|\Phi_{2}(x)\rangle=\frac{i}{2}\int_{{\cal R}^{3}}d^{3}x\rho(x)_{\Phi_{1}\Phi_{2}}\,,$ (3.31)
@@
$\rho(x)_{\Phi_{1}\Phi_{2}}=\Bigl{[}\Phi_{2}(x)\frac{\partial\Phi^{*}_{1}(x)}{\partial t}-\frac{\partial\Phi^{*}_{2}(x)}{\partial t}\Phi_{1}(x)\Bigr{]}\,,$ (3.32)
@@
$i\frac{\partial}{\partial t}\Psi=\hat{H}\Psi\,.$ (3.34)
@@
${\cal J}=\frac{i}{2}\Bigl{(}\Psi^{\dagger}\,\nabla\cdot\Psi-\Psi\,\nabla\cdot\Psi^{\dagger}\Bigr{)};\,\,\,\mbox{with}\,\,\nabla\cdot{\cal J}=0\,.$ (3.35)
@@
For Rovelli [44, 45, 46, 47] the absence of time is a feature of the classical Hamilton-Jacobi formulation of general relativity, and the wave function is only a function of the “3-geometry”, namely the equivalence class of metrics under a diffeomorphism, and not of the specific coordinate dependent form of the metric tensor. According to the second law of thermodynamics, forward in time represents the direction in which entropy increases and in which we obtain information, so the flow of time would represent a subjective feature of the universe, not an objective part of physical reality [44, 45, 46, 47]. In this realm, in which the observable universe does not exhibit time-reversal symmetry, events, rather than particles or fields, are the basic constituents of the universe, implying that the evolution of physical quantities is related to the description of the relationship between events [44, 45, 46, 47]. For instance, given the wave function of the universe as a functional constrained to a region configuration of a super-space that contains a three-surface $\Sigma$ and matter fields configuration, represented by $\phi$, where the metric is described by $h_{ij}$ the corresponding WdW wave function $\Psi(h_{ij},\phi)$ may be interpreted according to [44, 45, 46, 47], as stressed before, as describing the evolution of $\Psi(\phi)$ in the physical variable $\phi$. In the following we assume the proposal made by Hawking [57] for which $|\Psi[h_{ij},\phi,\Sigma]|^{2}$ is proportional to the probability $P({\cal A})$ of finding a 3-surface $\Sigma$ with metric $h_{ij}$ and matter field configuration $\phi$:
$*P({\cal A})\propto\int_{\cal A}|\Psi[h_{ij},\phi,\Sigma]|^{2}\,{}^{*}1$ (3.36)
@@
where $*{}^{*}$ is the Hoge dual in the metric $\Gamma(N)$, i.e. $*{}^{*}1$ is the volume element [57]. $\Gamma(N)$ in turn represents the superspace metric which does not depend linearly on $N$. In what follows, we adopt similar interpretations to Hawking’s prescription for $|\Psi(\eta)|^{2}$ and $|\Psi(\xi)|^{2}$, and to establish a connection with standard quantum mechanics we use the denomination “probability density".
@@
Briefly, as highlighted by Hartle [61], in standard quantum mechanics, the probabilities associated with wave functions are represented by squares of amplitudes, and additionally, a criterion is needed to specify which sets of histories can have probabilities consistently assigned to them. And yet, as highlighted by Hartle, in standard quantum mechanics that criterion is measurement – probabilities can be consistently assigned to histories of measured alternatives and usually not otherwise. Hartle then recalled that a criterion based on measurements or observers cannot be fundamental in a quantum theory that seeks to explain the early universe where neither existed. A more general criterion for closed systems assigns probabilities to just those sets of histories for which there is vanishing interference between its individual members as a consequence of the system’s initial quantum state. Such sets of histories are said to decohere. Decoherent sets of histories are what may be used for prediction and retrodiction in quantum cosmology for they may be assigned probabilities [61]. And this type of identification, in our view, with decoherent sets of histories, is in perfect harmony with Hawking’s prescription and with the modes of interpretation adopted by us for the dynamic evolutionary processes of the Universe’s wave function and the scale factor in branch-cut gravity.
3.2 The wave function of the Universe in the non-commutative approach
@@
In what follows, we use the powerful iterative Range-Kutta-Fehlberg numerical analysis for solving differential equations to find solutions for equation (3.23), without introducing numerical approximations, in order to describes the evolution of the wave function of the Universe, $\Psi(\eta)$. The boundary conditions of the solutions are based on the Bekenstein criterion, which provides an upper limit for the Universe’s entropy, following the proposition presented in [8, 9]. The total entropy of a black hole, according to the Bekenstein limit, is proportional to the number of Planck areas needed to cover the event horizon, where each area corresponds to one unit of entropy. In non-commutative branched gravitation, we assume that the primordial singularity is equally covered by a certain number of Planck areas, whose numerical value in turn corresponds to the total primordial entropy of the Universe. We assume that the dimensions of this boundary region correspond to the farthest points observable while still respecting causality. For this, we consider an appropriate distance, denoted as $d(t)$, between a pair of objects at any given moment $t$, and the corresponding distance, denoted as $d(t_{0})$, at a reference time $t_{0}$. We establish this relationship as $d(t)=|\eta(t)|d(t_{0})$. This means that the relationship between the two distances is modulated by the scale factor of the branch-cut Universe. This implies that for $t=t_{0}$ we have $|\eta(t_{0})|=1$. From a quantum probabilistic point of view, this condition implies a maximum probability of observation, $|\Psi(1)|^{2}=1$, assuming a normalized wave function. Thus, the boundary conditions considered in this contribution are, in the contraction sector $\Psi(-1)=-1$, while in the expansion region, $\Psi(1)=1$.
@@
Figures 1 and 2 depict the behavior of the solutions to equation (3.23), which pertains to the wave function of the Universe $\Psi(\eta)$ within a non-commutative approach. These results align with the predictions of branch-cut gravity , suggesting that the present Universe did not originate from nothing (see refs. [6, 7, 65]) or from a quantum loop [66]. Instead, it appears to have emerges from a prior phase before the current expansion phase. The implications of the non-commutative algebra on the acceleration of the Universe are most pronounced in figure 2, indicating the presence of a dynamic acceleration of the Universe driven by a force, whose work can be expressed as $W=-p_{f}dV$, where $p_{f}$ represents the strength of the pressure in the expanding region.
Figure 1: Typical sample solutions of the wave function of the Universe $\Psi(\eta)$ utilizing the non-commutative approach given by equation (3.23). The solutions are for initial conditions $\Psi(1)$ and $\Psi^{\prime}(1)$ and different values of the parameter $\gamma$. The parameter $\alpha$ was fixed at $1/3$. The parameter $\chi$ is implicitly embedded in the intrinsic structure of the scale factor $\eta$ as well as in the dual complementary scale factor $\xi$. The values of the running coupling constants and parameters are: $\tilde{g}_{r}=0.4$; $\tilde{g}_{m}=0.6185$; $g_{k}=1$; $g_{q}=0.7$; $g_{\Lambda}=0.333$; $g_{s}=0.03$; $\alpha=1/3$; ${\cal C}=1$.
@@
Figure 2: Typical solutions of the wave function of the Universe $\Psi(\eta)$ using the non-commutative approach given by equation (3.23). The parameter $\alpha$ was fixed at 1/3. The parameter $\chi$ is implicitly embedded in the intrinsic structure of the scale factor $\eta$ as well as in the dual complementary scale factor $\xi$. On the left side, in figures ((a), (c), and (e)) the values of the running coupling constants and parameters are as follows: $\tilde{g}_{r}=0.4$; $\tilde{g}_{m}=0.6185$; $g_{k}=1$; $g_{q}=0.7$; $g_{\Lambda}=0.333$; $g_{s}=-0.03$; $\alpha=1/3$; $\gamma=1$; ${\cal C}=1$. The boundary condition is set as $\Psi(-1)=-1$. On the right side, in figures ((b), (d), and (f)) , the values of the running coupling constants and parameters are: $\tilde{g}_{r}=0.4$; $\tilde{g}_{m}=0.6185$; $g_{k}=1$; $g_{q}=0.7$; $g_{\Lambda}=0.333$; $g_{s}=-0.03$; $\alpha=1/3$; $\gamma=-1$; ${\cal C}=1$. The boundary condition is set as $\Psi(1)=1$.
Another explanation could be based on the contribution of a new kind of dynamical energy fluid or field, that fills all space, whose effect on the expansion of the universe would be the opposite of that of “normal” matter and energy, as for instance the “quintessence" and “stiff” contributions, parameterized by the terms corresponding to the running coupling constants $g_{q}$ and $g_{s}$. Our results indicate that the term corresponding to stiff matter has expressive implications on the conformation of the potentials adopted in the present formulation, particularly with regard to the transition regions between the contraction and expansion phases of the branch-cut Universe.
Finally, when we compare the Universe expansion results corresponding to the commutative and non-commutative formulations, the configuration of matter, represented by the potentials (4.11), (4.15), (4.24), and (4.26), show expressive differences demonstrating that the non-commutative formulation imprints on the structure of spacetime a non-symmetrical redistribution of matter which captures, as stressed before, the short-distance scales.
@@
Figure 3: On the left side we show the complex plane solution of the wave function $\Psi(\xi)$, which is the dual complementary to the wave function of the Universe $\Psi(\eta)$ when utilizing the commutative approach given by equation (3.25). On the right side, we show the probability density of the wave function $\Psi(\xi)$ using Hawking’s prescription given by equation (3.36).
3.3 Dual spaces in quantum mechanics